# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ Appendix F Proof of an upper bound on the asymmetry cost
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil rN\rceil},\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime\otimes N}(\tilde{\sigma}_{N,g})\right)$ (427)
$=T\left(\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right)+p_{\mathrm{err}}\rho_{\mathrm{sym}},\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime}(\sigma_{\bm{i},g})\right)^{\otimes N}+p_{\mathrm{err}}\rho_{\mathrm{sym}}\right)$ (428)
$\leq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$ (429)
$\leq\max_{\bm{i}\in\mathcal{T}_{\delta}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right),$ (430)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)=0.$ (431)
@@ Appendix H QGT of purification
$O_{tot}\coloneqq O_{S}\otimes I_{A}+I_{S}\otimes O_{A},\quad O_{S}=\bm{\gamma}^{\dagger}\bm{X}_{S},\quad O_{A}=\bm{\gamma}^{\dagger}\bm{X}_{A}.$ (435)
We will minimize
@@
over purifications $\Psi_{\rho}$ of $\rho$ and the sets of Hermitian operators $\bm{X}_{A}$ on $A$. We remark that the QGT $\mathcal{Q}^{\Psi_{\rho}}$ of purification $\Psi_{\rho}$ satisfies $V\left(\Phi_{\rho},O_{tot}\right)=\bm{\gamma}^{\dagger}\mathcal{Q}^{\Psi_{\rho}}\bm{\gamma}$.
@@
$\braket{\Phi_{\rho}^{\prime}|O_{tot}(I-\Phi_{\rho})O_{tot}^{\dagger}|\Phi_{\rho}^{\prime}}=\braket{\Phi_{\rho}|O_{tot}^{\prime}(I-\Phi_{\rho})O_{tot}^{\prime{\dagger}}|\Phi_{\rho}},$ (437)
@@
for all $\mu$, implying that $\braket{\Phi_{\rho}|\bm{\gamma}^{\dagger}\bm{X}_{S}\otimes I_{A}+I_{S}\otimes\bm{\gamma}^{\dagger}\tilde{\bm{X}}_{A}|\Phi_{\rho}}=0$. Since adding the multiple of $I_{A}$ to $X_{A,\mu}$ does not change $V\left(\Phi_{\rho},O_{tot}\right)$, we assume, without loss of generality, that $\braket{\Phi_{\rho}|O_{tot}|\Phi_{\rho}}=\braket{\Phi_{\rho}|O_{tot}^{\dagger}|\Phi_{\rho}}=0$ in the following. In this case, we have
@@
Under a small variation of $\bm{X}_{A}$ to $\bm{X}_{A}+\delta\bm{X}_{A}$ for Hermitian operators $\delta\bm{X}_{A}$, the operator $O_{A}$ changes $O_{A}\to O_{A}+\delta O_{A}$, where $\delta O_{A}\coloneqq\bm{\gamma}^{\dagger}\delta\bm{X}_{A}$, and hence we get
@@
$=\sum_{\mu,\nu}\mathcal{K}_{\mu\nu}\gamma^{*\mu}\gamma^{\nu}+O((\delta O_{A})^{2}),$ (443)
@@
Since $\mathcal{K}$ is Hermitian, $\delta\left\|\mathrm{i}[\Phi_{\rho},O_{tot}]\right\|_{f_{q},\Phi_{\rho}}^{2}$ vanishes up to the first order with respect to $\delta O_{A}$ for all $\bm{\gamma}$ if and only if $\mathcal{K}=0$.
$\mathcal{K}_{\mu\mu}=0$ holds for all $\delta\bm{X}_{A}$ only if
@@
over purification $\Psi_{\rho}$ of $\rho$ and the set of Hermitian operators $\bm{X}_{A}$ on $A$. Note that $f_{q}$ for $q=1/2$ corresponds to the SLD norm. In this case, we find
@@
$=\sum_{\mu,\nu}\mathcal{M}_{\mu\nu}\gamma^{*\mu}\gamma^{\nu}+O((\delta O_{A})^{2}),$ (451)
@@
Since $\mathcal{M}$ is Hermitian, $\delta\left\|\mathrm{i}[\Phi_{\rho},O_{tot}]\right\|_{f_{q},\Phi_{\rho}}^{2}$ vanishes up to the first order with respect to $\delta O_{A}$ for all $\bm{\gamma}$ if and only if $\mathcal{M}=0$.
@@
which vanishes for $|X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}|X_{S,\mu}|\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}$ only for $q=1/2$. We remark that for $q=1/2$, $\mathcal{M}=0$ for the operators $X_{A,\mu}$ in Eq. (455), which corresponds to the result in [24] stating that the SLD quantum Fisher information satisfies $\mathcal{F}_{H}(\rho)=\min_{\Psi_{\rho},H_{A}}4V(H+H_{A},\Psi_{\rho})$, where the minimization is taken over all the purifications and the Hermitian operators $H_{A}$.
@@ Appendix I Properties of QGT and its extension
We here summarize the fundamental properties of the tensor $\mathcal{S}_{q}$ defined in Eq. (85). Taking the limit $q\to 1^{-}$, they also hold for $\mathcal{S}$, defined in Eq. (68), which coincides with the QGT $\mathcal{Q}$ for pure states. These properties are inherited from those of Petz’s monotone metric:
Proposition 43 (Properties of $\mathcal{S}_{q}$).
@@
    Positivity: For any state $\rho$, $\mathcal{S}_{q}^{\rho}\geq 0$.
@@
    Additivity: For any pair of states $\rho$ and $\sigma$, $\mathcal{S}_{q}^{\rho\otimes\sigma}=\mathcal{S}_{q}^{\rho}+\mathcal{S}_{q}^{\sigma}$.
@@
    Convexity: For any probability distribution $\{p_{k}\}_{k}$ and set of states $\{\rho_{k}\}_{k}$, $\sum_{k}p_{k}\mathcal{S}_{q}^{\rho_{k}}\geq\mathcal{S}_{q}^{\sum_{k}p_{k}\rho_{k}}$
@@
    Monotonicity: For any $G$-covariant quantum channel $\mathcal{E}$ and state $\rho$, $\mathcal{S}_{q}^{\rho}\geq\mathcal{S}_{q}^{\mathcal{E}(\rho)}$.
@@
$Y_{\mu}\coloneqq\frac{\partial}{\partial\lambda^{\mu}}V(g(\bm{\lambda}))\biggl{|}_{\bm{\lambda}=\bm{0}}=X_{\mu}\otimes\mathbb{I},$ (461)
where $X_{\mu}\coloneqq\frac{\partial}{\partial\lambda^{\mu}}U(g(\bm{\lambda}))|_{\bm{\lambda}=\bm{0}}$. By applying Eq. (458) for $O\coloneqq\bm{\gamma}^{\dagger}\bm{X}$ with $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we get
$\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\bm{\gamma}^{\dagger}\bm{X}]\|_{f,\rho_{k}}^{2}=\left\|\mathrm{i}\left[\sigma,\bm{\gamma}^{\dagger}\bm{Y}\right]\right\|_{f,\sigma}^{2}.$ (462)
@@
$\left\|\mathrm{i}\left[\sigma,\bm{\gamma}^{\dagger}\bm{Y}\right]\right\|_{f_{q},\sigma}^{2}\geq\left\|\mathrm{i}\left[\sum_{k}p_{k}\rho_{k},\bm{\gamma}^{\dagger}\bm{X}\right]\right\|_{f_{q},\sum_{k}p_{k}\rho_{k}}^{2}.$ (463)
@@
$\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\bm{\gamma}^{\dagger}\bm{X}]\|_{f_{q},\rho_{k}}^{2}\geq\left\|\mathrm{i}\left[\sum_{k}p_{k}\rho_{k},\bm{\gamma}^{\dagger}\bm{X}\right]\right\|_{f_{q},\sum_{k}p_{k}\rho_{k}}^{2}.$ (464)
@@
$\sum_{k}p_{k}\bm{\gamma}^{\dagger}\mathcal{S}_{q}^{\rho_{k}}\bm{\gamma}\geq\bm{\gamma}^{\dagger}\mathcal{S}_{q}^{\sum_{k}p_{k}\rho_{k}}\bm{\gamma}$ (465)
holds for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, which completes the proof of the property (iii).
@@
$\left\|\mathrm{i}\left[\rho,\bm{\gamma}^{\dagger}\bm{X}\right]\right\|_{f_{q},\rho}^{2}\geq\left\|\mathrm{i}\left[\sigma,\bm{\gamma}^{\dagger}\bm{Y}\right]\right\|_{f_{q},\sigma}^{2}.$ (467)
@@
$\left\|\mathrm{i}\left[\rho,\bm{\gamma}^{\dagger}\bm{X}\right]\right\|_{f_{q},\rho}^{2}\geq\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\bm{\gamma}^{\dagger}\bm{X}]\|_{f_{q},\rho_{k}}^{2}.$ (468)
@@
$\bm{\gamma}^{\dagger}\mathcal{S}_{q}^{\rho}\bm{\gamma}\geq\sum_{k}p_{k}\bm{\gamma}^{\dagger}\mathcal{S}_{q}^{\rho_{k}}\bm{\gamma}$ (469)
holds for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, which completes the proof of the property (v). ∎

### v4

@@ Appendix F Proof of an upper bound on the asymmetry cost (Proposition 7)
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil rN\rceil},\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\prime\otimes N}(\tilde{\sigma}_{N,g})\right)$ (453)
$=T\left(\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right)+p_{\mathrm{err}}\rho_{\mathrm{sym}},\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\prime}(\sigma_{\bm{i},g})\right)^{\otimes N}+p_{\mathrm{err}}\rho_{\mathrm{sym}}\right)$ (454)
$\leq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$ (455)
$\leq\max_{\bm{i}\in\mathcal{T}_{\delta}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right),$ (456)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)=0.$ (457)
@@ Appendix H QGT of purification
$O_{tot}\coloneqq O_{S}\otimes I_{A}+I_{S}\otimes O_{A},\quad O_{S}=\gamma^{\dagger}X_{S}\coloneqq\gamma^{\mu*}X_{S,\mu},\quad O_{A}=\gamma^{\dagger}X_{A}\coloneqq\gamma^{\mu*}X_{A,\mu}$ (461)
for $\gamma\in\mathbb{C}^{\dim G}$. We introduce $|\Phi_{\rho}\coloneqq\ket{\Phi_{\rho}}\bra{\Phi_{\rho}}$. We will minimize
@@
over purifications $\Phi_{\rho}$ of $\rho$ and the sets of Hermitian operators $\{X_{A,\mu}\}_{\mu=1}^{\dim G}$ on $A$. We remark that the QGT $\mathcal{Q}^{\Phi_{\rho}}$ of purification $\Phi_{\rho}$ satisfies $V\left(\Phi_{\rho},O_{tot}\right)=\gamma^{\dagger}\mathcal{Q}^{\Phi_{\rho}}\gamma$.
@@
$\braket{\Phi_{\rho}^{\prime}|O_{tot}(I-\Phi_{\rho}^{\prime})O_{tot}^{\dagger}|\Phi_{\rho}^{\prime}}=\braket{\Phi_{\rho}|O_{tot}^{\prime}(I-\Phi_{\rho})O_{tot}^{\prime{\dagger}}|\Phi_{\rho}},$ (463)
@@
for all $\mu$, implying that $\braket{\Phi_{\rho}|\gamma^{\dagger}X_{S}\otimes I_{A}+I_{S}\otimes\gamma^{\dagger}\tilde{X}_{A}|\Phi_{\rho}}=0$. Since adding the multiple of $I_{A}$ to $X_{A,\mu}$ does not change $V\left(\Phi_{\rho},O_{tot}\right)$, we assume, without loss of generality, that $\braket{\Phi_{\rho}|O_{tot}|\Phi_{\rho}}=\braket{\Phi_{\rho}|O_{tot}^{\dagger}|\Phi_{\rho}}=0$ in the following. In this case, we have
@@
Under a small variation of $X_{A,\mu}$ to $X_{A,\mu}+\delta X_{A,\mu}$ for $\mu=1,\ldots,\dim G$ with Hermitian operators $\delta X_{A,\mu}$, the operator $O_{A}$ changes $O_{A}\to O_{A}+\delta O_{A}$, where $\delta O_{A}\coloneqq\gamma^{\mu*}\delta X_{A,\mu}$, and hence we get
@@
$=\mathcal{K}_{\mu\nu}\gamma^{\mu*}\gamma^{\nu}+O((\delta O_{A})^{2}),$ (469)
@@
Since $\mathcal{K}$ is Hermitian, $\delta\left\|\mathrm{i}[\Phi_{\rho},O_{tot}]\right\|_{f_{q},\Phi_{\rho}}^{2}$ vanishes up to the first order with respect to $\delta O_{A}$ for all $\gamma$ if and only if $\mathcal{K}=0$.
$\mathcal{K}_{\mu\mu}=0$ holds for all $\{\delta X_{A,\mu}\}_{\mu=1}^{\dim G}$ only if
@@
over purification $\Phi_{\rho}$ of $\rho$ and the set of Hermitian operators $\{X_{A,\mu}\}_{\mu=1}^{\dim G}$ on $A$. Note that $f_{q}$ for $q=1/2$ corresponds to the SLD norm. In this case, we find
@@
$=\mathcal{M}_{\mu\nu}\gamma^{\mu*}\gamma^{\nu}+O((\delta O_{A})^{2}),$ (477)
@@
Since $\mathcal{M}$ is Hermitian, $\delta\left\|\mathrm{i}[\Phi_{\rho},O_{tot}]\right\|_{f_{q},\Phi_{\rho}}^{2}$ vanishes up to the first order with respect to $\delta O_{A}$ for all $\gamma$ if and only if $\mathcal{M}=0$.
@@
which vanishes for $|X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}|X_{S,\mu}|\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}$ only for $q=1/2$. We remark that for $q=1/2$, $\mathcal{M}=0$ for the operators $X_{A,\mu}$ in Eq. (481), which corresponds to the result in [24] stating that the SLD quantum Fisher information satisfies $\mathcal{F}_{H}(\rho)=\min_{\Phi_{\rho},H_{A}}4V(H+H_{A},\Phi_{\rho})$, where the minimization is taken over all the purifications and the Hermitian operators $H_{A}$.
@@ Appendix I Properties of QGT and its extension
We here summarize the fundamental properties of the tensor $\mathcal{Q}_{q}$ defined in Eq. (77). Taking the limit $q\to 1^{-}$, they also hold for $\mathcal{Q}$, defined in Eq. (43), which coincides with the QGT $\mathcal{Q}$ for pure states. These properties are inherited from those of Petz’s monotone metric:
Proposition 43 (Properties of $\mathcal{Q}_{q}$).
@@
    Positivity: For any state $\rho$, $\mathcal{Q}_{q}^{\rho}\geq 0$.
@@
    Additivity: For any pair of states $\rho$ and $\sigma$, $\mathcal{Q}_{q}^{\rho\otimes\sigma}=\mathcal{Q}_{q}^{\rho}+\mathcal{Q}_{q}^{\sigma}$.
@@
    Convexity: For any probability distribution $\{p_{k}\}_{k}$ and set of states $\{\rho_{k}\}_{k}$, $\sum_{k}p_{k}\mathcal{Q}_{q}^{\rho_{k}}\geq\mathcal{Q}_{q}^{\sum_{k}p_{k}\rho_{k}}$
@@
    Monotonicity: For any $G$-covariant quantum channel $\mathcal{E}$ and state $\rho$, $\mathcal{Q}_{q}^{\rho}\geq\mathcal{Q}_{q}^{\mathcal{E}(\rho)}$.
@@
$Y_{\mu}\coloneqq\frac{\partial}{\partial\lambda^{\mu}}V(g(\lambda))\biggl|_{\lambda=0}=X_{\mu}\otimes\mathbb{I},$ (487)
where $X_{\mu}\coloneqq\frac{\partial}{\partial\lambda^{\mu}}U(g(\lambda))|_{\lambda=0}$. By applying Eq. (484) for $O\coloneqq\gamma^{\dagger}X$ with $\gamma\in\mathbb{C}^{\dim G}$, we get
$\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\gamma^{\dagger}X]\|_{f,\rho_{k}}^{2}=\left\|\mathrm{i}\left[\sigma,\gamma^{\dagger}Y\right]\right\|_{f,\sigma}^{2}.$ (488)
@@
$\left\|\mathrm{i}\left[\sigma,\gamma^{\dagger}Y\right]\right\|_{f_{q},\sigma}^{2}\geq\left\|\mathrm{i}\left[\sum_{k}p_{k}\rho_{k},\gamma^{\dagger}X\right]\right\|_{f_{q},\sum_{k}p_{k}\rho_{k}}^{2}.$ (489)
@@
$\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\gamma^{\dagger}X]\|_{f_{q},\rho_{k}}^{2}\geq\left\|\mathrm{i}\left[\sum_{k}p_{k}\rho_{k},\gamma^{\dagger}X\right]\right\|_{f_{q},\sum_{k}p_{k}\rho_{k}}^{2}.$ (490)
@@
$\sum_{k}p_{k}\gamma^{\dagger}\mathcal{Q}_{q}^{\rho_{k}}\gamma\geq\gamma^{\dagger}\mathcal{Q}_{q}^{\sum_{k}p_{k}\rho_{k}}\gamma$ (491)
holds for any $\gamma\in\mathbb{C}^{\dim G}$, which completes the proof of the property (iii).
@@
$\left\|\mathrm{i}\left[\rho,\gamma^{\dagger}X\right]\right\|_{f_{q},\rho}^{2}\geq\left\|\mathrm{i}\left[\sigma,\gamma^{\dagger}Y\right]\right\|_{f_{q},\sigma}^{2}.$ (493)
@@
$\left\|\mathrm{i}\left[\rho,\gamma^{\dagger}X\right]\right\|_{f_{q},\rho}^{2}\geq\sum_{k}p_{k}\|\mathrm{i}[\rho_{k},\gamma^{\dagger}X]\|_{f_{q},\rho_{k}}^{2}.$ (494)
@@
$\gamma^{\dagger}\mathcal{Q}_{q}^{\rho}\gamma\geq\sum_{k}p_{k}\gamma^{\dagger}\mathcal{Q}_{q}^{\rho_{k}}\gamma$ (495)
holds for any $\gamma\in\mathbb{C}^{\dim G}$, which completes the proof of the property (v). ∎
@@
Appendix K Circumventing constraint imposed by symmetry subgroup
When converting a pure state $\psi$ to another pure state $\phi$ in the asymptotic setup, the monotonicity of the QGTs yields an upper bound on possible conversion rate as shown in Sec. IV.1. However, the symmetry subgroups further restrict achievable conversion rate: the conversion rate always vanishes if $\mathrm{Sym}_{G}(\psi)\not\subset\mathrm{Sym}_{G}(\phi)$. In this section, we investigate ways to circumvent this restriction due to the symmetry subgroups by modifying the setup.
K.1 Synchronization of asymmetry
$=\frac{V(\psi,H)}{V(\phi,H^{\prime})}+1,$ (516)
regardless of the periods of $\psi$ and $\phi$. Importantly, even when $R(\psi\to\phi)=0$, $\{\psi^{\otimes N}\otimes\phi^{\otimes N}\}_{N}$ is asymptotically convertible to $\{\phi^{\otimes\lfloor Nr^{*}\rfloor}\otimes\phi^{\otimes N}\}_{N}$ for any $r^{*}\in[0,V(\psi,H)/V(\phi,H^{\prime}))$. In this case, therefore, $\phi^{\otimes N}$ acts like a catalyst, enhancing the conversion rate from zero to $V(\psi,H)/V(\phi,H^{\prime})$ by bypassing the restriction on state convertibility imposed by their periods.
$R(\psi\otimes\phi\to\phi)$
K.2 Sublinear additional resource
Theorem 46.
See Fig. 6 for a schematic picture of the setup of Theorem 46.
Proof of Theorem 46.
$\quad\geq\frac{M}{N}\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma-\frac{M}{N}h(\epsilon)+\frac{1}{N}o\left(M\right),$ (520)
where $O^{\prime\prime}$ is defined by $O^{\prime\prime}\coloneqq\gamma^{\dagger}X^{\prime\prime}$ and $X^{\prime\prime}_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U^{\prime\prime}(g(\lambda))|_{\lambda=0}$ by using the projective unitary representation of $G$ on the Hilbert space on which $\chi$ is defined. Since $\|\mathrm{i}[\chi,O^{\prime\prime}]\|_{f_{q},\chi}^{2}<\infty$ for any finite-dimensional system, the second term in the left-hand side vanishes in the limit of $N\to\infty$. By following the arguments from Eq. (35) to Eq. (37), we therefore get
Direct part: In the estimation step, we estimate $g\in G$ by consuming $(\mathcal{U}_{g}(\psi)\otimes\mathcal{U}_{g}^{\prime\prime}(\chi))^{\otimes\lceil N^{1-\epsilon}\rceil}$. The estimator $\hat{g}$ ensured to exist by Lemma 26 satisfies $((\mathcal{U}_{\theta}\otimes\mathcal{U}^{\prime\prime}_{\theta})\circ(\mathcal{U}_{\hat{g}}\otimes\mathcal{U}_{\hat{g}}^{\prime\prime}))(\psi\otimes\chi)=(\mathcal{U}_{g}\otimes\mathcal{U}^{\prime\prime}_{g})(\psi\otimes\chi)$, i.e.,
$\mathcal{U}_{\theta}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)\,\land\,\mathcal{U}^{\prime\prime}_{\theta}\circ\mathcal{U}^{\prime\prime}_{\hat{g}}(\chi)=\mathcal{U}^{\prime\prime}_{g}(\chi)$ (522)
with a failure probability that asymptotically vanishes as $N\to\infty$. From the assumption $\mathrm{Sym}_{G}(\chi)\subset\mathrm{Sym}_{G}(\phi)$, if $\theta$ satisfies $\mathcal{U}^{\prime\prime}_{\theta}\circ\mathcal{U}^{\prime\prime}_{\hat{g}}(\chi)=\mathcal{U}^{\prime\prime}_{g}(\chi)$, then $\mathcal{U}^{\prime}_{\theta}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}^{\prime}_{g}(\phi)$ also holds, implying that Eq. (236) remains valid in this case.
We slightly modify the channel in Eq.(235) as